# Chenxu Wang

### v4

@@ Abstract
The accurate prediction of drug molecule solubility is essential for determining their therapeutic effectiveness and safety, influencing the drug’s ADME processes. Traditional solubility prediction techniques often fail to capture the complex nature of molecular structures, leading to notable deviations between predictions and actual results. For example, the Discussion on Advanced Drug-Like Compound Structures. Lusci[1] highlighted issues in capturing crucial cyclic structural information in molecules with ring structures. To overcome this issue, our research introduces a novel deep learning framework combining attention-based transformers, Long Short-Term Memory (LSTM) networks, and Graph Convolutional Networks (GCN), aimed at enhancing the precision of solubility predictions. Utilizing a training set of 9,943 compounds and testing on an anticancer compound dataset, our method achieved a correlation coefficient ($R^{2}$) of 0.59 and a Root Mean Square Error (RMSE) of 0.57, which outperforms the benchmark models’ scores of 0.52 ($R^{2}$) and 0.61 (RMSE). Importantly, in an additional independent test, our model significantly outperformed the baseline with an RMSE of 1.05 compared to 1.28, a relative accuracy improvement of 45.9%. This research not only demonstrates the vast potential of deep learning for improving solubility prediction accuracy but also offers novel insights for drug design and selection in the future. Continued efforts will be directed towards optimizing the model architecture and extending its application to better support the drug development process, underscoring the pivotal role of deep learning in drug discovery.
@@ 1 Introduction
In contemporary medical science, the development and deployment of organic pharmaceuticals hold immeasurable value for public health improvement[2]. Organic drugs have emerged as heroes in the fight against chronic illnesses such as cancer, cardiovascular diseases, and diabetes. For example, some targeted therapies have become critical components of cancer management, markedly enhancing patient survival[3]. However, despite advancements in technology facilitating new therapeutic approaches, the path of drug development is fraught with challenges and high costs. In the U.S., the cost of bringing a new drug to market ranges from \$157.3 million to \$1.9508 billion and spans twelve to fifteen years from initial research to commercial availability[4, 5]. In this context, the precise prediction of a drug molecule’s water solubility is especially vital, as it directly influences the drug’s ADME properties, crucial for ensuring both its efficacy and safety[6].
While traditional methods for predicting water solubility—such as empirical models (which depend on pre-existing solubility data like the Jouyban-Acree model), quantitative structure-activity relationships (QSAR) models (building mathematical models based on the analysis of physicochemical properties and biological activities of compounds), and physicochemical computational methods (utilizing simulations of molecular interactions and thermodynamic properties through computational approaches like molecular dynamics and quantum mechanics)—have shown some efficacy, as demonstrated by Munjal’s QSAR model for Paclitaxel derivatives, they frequently underperform when applied to molecules with complex structures[7, 8, 9]. Compared to machine learning techniques, the primary limitations of these traditional methods include:
  1. 1.
    Restricted predictive accuracy, notably for molecules that are structurally complex or exhibit novel characteristics[10];
  2. 2.
    A heavy reliance on extensive experimental data, limiting effectiveness in cases with few samples or novel molecular structures[11];
  3. 3.
    The requirement for expensive computational resources, especially in physicochemical approaches[12];
  4. 4.
    Insufficient comprehension of intricate molecular interactions, hindering a comprehensive reflection of molecular solubility behavior[13].
These challenges have prompted the scientific community to explore more advanced and precise methods of prediction.
Within the current pharmaceutical research landscape, deep learning provides a novel approach that significantly improves the precision and efficiency of solubility predictions. Deep learning offers distinct potential advantages over traditional methodologies:
  1. 1.
    Automatic Feature Learning: Deep learning models inherently learn and extract valuable features from complex data without relying on predefined physicochemical parameters, enabling the handling of highly complex molecular structures[14].
  2. 2.
    Capacity to Handle Large Datasets: As scientific data proliferates, deep learning adeptly manages and analyzes extensive datasets, including those derived from high-throughput experiments, which boosts the predictive model’s generalizability and accuracy[15].
  3. 3.
    Adaptability and Flexibility of Models: Deep learning models can be tailored to predict particular molecular types by modifying network structures and parameters, such as employing various neural network types like CNNs and RNNs to better process spatial and temporal molecular data[16].
  4. 4.
    Swift Iteration and Updating: Deep learning models rapidly integrate new experimental data and align with emerging research and market needs, thereby expediting the drug development timeline[17].
Therefore, deep learning is chosen as the primary predictive model due to its outstanding performance and strong adaptability to unknown or complex data, significantly expediting the drug development process and aiding in the early identification of candidates with high therapeutic potential, thus greatly reducing both cost and time in development.
@@
Numerous studies have illustrated the role of deep learning in the fields of drug discovery and materials science. For instance, Mater et al. discussed deep learning’s applications in drug and material design and synthesis planning, underscoring its potential to enhance and expedite the drug development workflow[18]. Aliper et al. showed how deep neural networks could categorize drugs using transcriptome data, which helps predict drug effectiveness and side effects more accurately[19]. Additionally, Walters et al. applied deep learning to predict drug molecule properties, offering essential insights during the initial phases of drug design to circumvent expensive modifications during later development stages[20]. These investigations not only highlight deep learning’s impact in pharmaceutical research but also its effectiveness in tackling complex drug design challenges. Furthermore, Korotcov analyzed the performance and advantages of deep learning over other machine learning techniques in various drug discovery datasets[21]. Aliper explored predicting drugs’ pharmacological features using deep learning and discussed the prospects for drug repurposing[19]. Tang highlighted the use of deep learning in drug discovery, focusing on its potential to forecast drug-target interactions, devise new medications, and anticipate toxicities and side effects[22]. Sarode summarized deep learning techniques and their applications in drug design, including CNNs, RNNs, VAEs, and GANs[23]. Lastly, Qian et al. introduced a deep learning-based novel drug design system optimized for scenarios with limited data, utilizing one-shot and reinforcement learning to operate effectively in data-scarce environments[24].
@@
Beyond applications in drug and materials discovery, deep learning has been widely utilized for predicting solubility in small drug molecules. For instance, Alessandro employed a novel approach using recurrent neural networks to tackle the loss of cyclic structural information[1]. Recurrent neural networks not only address this issue but also enable the automatic learning of appropriate molecular representations, decreasing the need for extensive domain expertise. Not limited to simple RNNs, Waqar added attention mechanisms to convolutional neural networks, facilitating the broad, automatic acquisition of solubility-relevant chemical features[25]. In addition to neural network architectural innovations, Tong suggested using molecular descriptors instead of structural forms for training neural networks[26]. This approach not only simplifies the representation of complex molecular structures but also improves training efficiency. Moreover, researchers like Sumin have made innovative improvements to molecular inputs, with the MFPCP method incorporating molecular fingerprints and physicochemical properties into a combined descriptor that allows for high-performance solubility prediction[27].
In particular, to tackle unresolved challenges, this research explores and integrates cutting-edge deep learning models, ultimately innovatively utilizing a combination of Graph Convolutional Networks[28] (GCN) and Long Short-Term Memory networks[29] (LSTM), enhanced by a Transformer architecture to develop the YZS-Model. This model, which comprehends and analyzes SMILES strings, outputs predictions about the solubility of specific compounds. Our model adeptly merges multiple deep learning technologies to capture both the overall graph structure and sequential information of molecules, introducing an innovative method for predicting the solubility of drugs. The approach begins with the use of a Graph Convolutional Network (GCN) to capture the holistic graph structure of molecules, facilitating an understanding of their complex interconnections. Following this, Long Short-Term Memory networks (LSTM) and Transformer are employed to handle the sequential data within molecules. The Transformer’s self-attention mechanism is particularly effective in recognizing long-distance molecular dependencies, complemented by the LSTM’s ability to process and integrate aspects of molecular relationships not entirely captured by the Transformer[30].
@@
This multifaceted approach, which leverages various deep learning techniques, significantly enhances the model’s ability to understand and predict molecular properties. Specifically, the use of GCN allows for a comprehensive analysis of molecular topology, capturing key structural features. The Transformer, with its robust self-attention mechanism, efficiently identifies and processes long-range dependencies within molecules. The LSTM complements the Transformer, ensuring that the model fully understands and integrates complex sequence information. This multi-technique methodology provides a fresh perspective for predicting drug solubility, demonstrating the vast potential of deep learning in chemical data analysis and pharmaceutical development.
@@ 2.2 Data Preprocessing
By employing the RDKit library, SMILES strings can be converted into molecular objects, which serve as a foundation for further extraction and analysis of molecular features. For this purpose, we constructed a feature tensor to store detailed information about each molecule’s core attributes. These features include the type of atoms, the number of bonds between atoms, formal charges on atoms, the number of free radical electrons, hybridization types, whether the atom is part of an aromatic system, the number of directly bonded hydrogen atoms, and critical information regarding chirality and stereochemistry, as detailed in Table 1. Most attributes, such as the type of elements, are represented using binary encoding—where 0 indicates absence and 1 indicates presence. A few features, like the number of directly bonded hydrogen atoms, are represented using one-hot encoding.
These details are subsequently integrated into a ‘torch_geometric’ Data object, which encompasses the node features, the connectivity of edges, and their attributes, thereby forming a complete graph data structure.
@@
In our research, we adopted the version of the Cui2020 dataset segmented by Ahmad et al. for the training and testing phases of our machine learning model[25, 14]. The team led by Ahmad et al. randomly split the original Cui2020 dataset into ten equal and independent subsets, each termed a "fold." To evaluate our model’s efficacy, we implemented a 10-fold cross-validation technique: during each cycle, the model trained on nine folds and was tested on the remaining one. This procedure was replicated ten times, ensuring that each fold was used as a test set once.
@@ 2.9 YZS-Model
The model begins by employing a GCN to process the graphical representation of drug molecules, effectively unveiling their spatial structure. The GCN layer abstracts intricate feature vectors that describe molecular topologies by analyzing interactions between molecular atoms and their chemical bonds, providing a robust foundation for detailed structural analysis.
The model then utilizes the Transformer architecture to enhance the feature output from the GCN layer. Utilizing self-attention, the Transformer discerns broad dependencies among atoms within the molecule, capturing both microstructural details and comprehensive internal interactions, thus offering an advanced representation enriched with global information.
The LSTM layer further refines the understanding of features produced by the Transformer, capturing temporal dependencies inherent in the molecular structure sequences. This step is pivotal for uncovering the dynamic nature of molecular features across sequences, thereby enriching the solubility predictions with temporal insights.
Ultimately, the model’s fully connected layers transform the LSTM outputs into final solubility predictions, synthesizing high-dimensional features from the GCN, Transformer, and LSTM, and associating these features with the drug molecules’ solubility through nonlinear transformations. The design of the entire model is focused on effectively translating the spatial and temporal characteristics of molecules into their solubility properties, thus ensuring high precision and reliability in the solubility predictions for organic drugs.

### v5

@@ Abstract
Accurate prediction of drug molecule solubility is crucial for therapeutic effectiveness and safety. Traditional methods often miss complex molecular structures, leading to inaccuracies. We introduce the YZS-Model, a deep learning framework that innovatively integrates Graph Convolutional Networks (GCN), Transformer architectures, and Long Short-Term Memory (LSTM) networks to enhance prediction precision.GCNs excel at capturing intricate molecular topologies by modeling the relationships between atoms and bonds, providing a detailed understanding of the molecular structure. Transformers, with their self-attention mechanisms, effectively identify long-range dependencies within molecules, capturing global interactions that other models might overlook. LSTMs contribute by processing sequential data, preserving long-term dependencies and integrating temporal information within molecular sequences. This multifaceted approach leverages the strengths of each component, resulting in a model that comprehensively understands and predicts molecular properties.Trained on 9,943 compounds and tested on an anticancer dataset, the YZS-Model achieved an $R^{2}$ of 0.59 and an RMSE of 0.57, outperforming benchmark models 0.52($R^{2}$) and 0.61(RMSE). In an independent test, it demonstrated an RMSE of 1.05, improving accuracy by 45.9%.The integration of these deep learning techniques allows the YZS-Model to automatically learn valuable features from complex data without predefined parameters, efficiently handle large datasets, and adapt to various molecular types. This comprehensive capability significantly improves predictive accuracy and model generalizability. Its precision in solubility predictions can expedite drug development by optimizing candidate selection, reducing costs, and enhancing efficiency. Our research underscores deep learning’s transformative potential in pharmaceutical science, particularly for solubility prediction and drug design.
@@ 1 Introduction
In contemporary medical science, the development and application of organic pharmaceuticals hold immeasurable value for improving public health[1]. Organic drugs have become critical in combating chronic diseases such as cancer, cardiovascular diseases, and diabetes. For example, some targeted therapies have become essential components of cancer management, significantly improving patient survival rates[2]. However, despite technological advancements leading to new therapeutic approaches, the path of drug development remains fraught with challenges and high costs. In the United States, the cost of bringing a new drug to market ranges from \$157.3 million to \$1.9508 billion, spanning twelve to fifteen years from initial research to commercial availability[3, 4]. This underscores the necessity for efficient and cost-effective methods to streamline the drug development process.
One of the critical aspects of drug development is the accurate prediction of a drug molecule’s water solubility, as it directly influences the drug’s ADME (Absorption, Distribution, Metabolism, and Excretion) properties, ensuring its efficacy and safety[5]. Accurate solubility predictions can significantly reduce the risk of late-stage failures and optimize the drug design process. Traditional methods for predicting water solubility, such as empirical models (e.g., Jouyban-Acree model), quantitative structure-activity relationship (QSAR) models, and physicochemical computational methods, although effective in some cases, often perform poorly when applied to molecules with complex structures[6, 7, 8]. These methods have the following main limitations: restricted predictive accuracy, especially for molecules that are structurally complex or have novel characteristics[9]; a heavy reliance on extensive experimental data, limiting their effectiveness in cases with few samples or novel molecular structures[10]; the requirement for expensive computational resources, particularly in physicochemical approaches[11]; and insufficient comprehension of intricate molecular interactions, hindering a comprehensive reflection of molecular solubility behavior[12]. These challenges have prompted the scientific community to explore more advanced and precise prediction methods.
Within the current pharmaceutical research landscape, deep learning provides a novel approach that significantly improves the precision and efficiency of solubility predictions. Deep learning offers distinct potential advantages over traditional methodologies: automatic feature learning, where deep learning models can inherently learn and extract valuable features from complex data without relying on predefined physicochemical parameters, enabling the handling of highly complex molecular structures[13]; the capacity to handle large datasets, as deep learning can adeptly manage and analyze extensive datasets, including those derived from high-throughput experiments, enhancing the predictive model’s generalizability and accuracy[14]; adaptability and flexibility of models, allowing deep learning models to be tailored to predict particular molecular types by modifying network structures and parameters, such as employing various neural network types like CNNs and RNNs to better process spatial and temporal molecular data[15]; and swift iteration and updating, enabling deep learning models to rapidly integrate new experimental data and align with emerging research and market needs, thereby expediting the drug development timeline[16]. Therefore, deep learning is chosen as the primary predictive model due to its outstanding performance and strong adaptability to unknown or complex data, significantly expediting the drug development process and aiding in the early identification of candidates with high therapeutic potential, thus greatly reducing both cost and time in development.
@@
Numerous studies have demonstrated the significant role of deep learning in the fields of drug discovery and materials science. For example, Mater et al. discussed the applications of deep learning in drug and material design and synthesis planning, highlighting its potential to enhance and expedite the drug development workflow[17]. Aliper et al. showed how deep neural networks can categorize drugs using transcriptome data, which helps predict drug effectiveness and side effects more accurately[18]. Additionally, Walters et al. applied deep learning to predict drug molecule properties, providing essential insights during the initial phases of drug design to circumvent expensive modifications during later development stages[19].
@@
Deep learning has also demonstrated its importance and effectiveness in predicting the solubility of small molecule drugs. For instance, Alessandro employed a novel approach using recurrent neural networks to tackle the issue of cyclic structural information loss[20]. Recurrent neural networks not only address this issue but also enable the automatic learning of appropriate molecular representations, reducing the need for extensive domain expertise. Beyond simple RNNs, Waqar added attention mechanisms to convolutional neural networks, facilitating the broad and automatic acquisition of solubility-relevant chemical features[21]. In addition to innovations in neural network architecture, Tong suggested using molecular descriptors instead of structural forms for training neural networks[22]. This approach not only simplifies the representation of complex molecular structures but also improves training efficiency. Moreover, researchers such as Sumin have made innovative improvements to molecular inputs, with the MFPCP method incorporating molecular fingerprints and physicochemical properties into a combined descriptor for high-performance solubility prediction[23]. These studies not only emphasize the impact of deep learning in pharmaceutical research but also demonstrate its effectiveness in addressing complex drug design challenges.
To address the unresolved challenges in predicting drug solubility, our research innovatively integrates advanced deep learning models to develop the YZS-Model. Traditional methods for predicting water solubility—such as empirical models, QSAR models, and physicochemical computational methods—have shown limited success, particularly with structurally complex or novel molecules. These methods often suffer from restricted predictive accuracy, a heavy reliance on extensive experimental data, high computational costs, and an inadequate understanding of intricate molecular interactions. The YZS-Model addresses these critical issues by combining Graph Convolutional Networks[24] (GCNs), Long Short-Term Memory networks[25] (LSTM), and Transformer[26] architecture to analyze SMILES strings and predict compound solubility. This innovative model leverages multiple deep learning technologies to capture both the global graph structure and sequential information of molecules, representing a novel approach to drug solubility prediction.
@@
The methodology begins with GCNs to capture the holistic graph structure of molecules, enabling a detailed understanding of their complex interconnections. Following this, LSTM and Transformer networks handle the sequential data within molecules. The Transformer’s self-attention mechanism is particularly effective at recognizing long-distance molecular dependencies, while the LSTM processes and integrates molecular relationships not entirely captured by the Transformer. This multifaceted approach significantly enhances the model’s ability to understand and predict molecular properties.Specifically, the GCN component allows for a comprehensive analysis of molecular topology, capturing key structural features. The Transformer’s robust self-attention mechanism efficiently identifies and processes long-range dependencies within molecules[26]. Meanwhile, the LSTM ensures that the model fully understands and integrates complex sequence information. This multi-technique methodology provides a fresh perspective for predicting drug solubility, demonstrating the vast potential of deep learning in chemical data analysis and pharmaceutical development.
Through these innovations, the YZS-Model not only addresses the limitations of traditional methods but also significantly reduces the risk of late-stage drug development failures. Accurate solubility predictions can optimize the drug design process, leading to more effective and safer drugs. Therefore, this research is essential for advancing pharmaceutical sciences, improving public health outcomes, and reducing the overall cost and time associated with bringing new drugs to market.
@@
The main contributions of this paper are as follows:
Multi-model Integration and Interaction Design: The core innovation of the YZS-Model lies in its unique multi-model integration and interaction design, combining GCNs, LSTM networks, and Transformer architecture. This design enables the model to comprehensively capture both local and global information of molecules, significantly enhancing the accuracy and robustness of predicting organic drug solubility. By integrating multiple advanced deep learning models, the YZS-Model successfully achieves a multi-level and multi-perspective analysis of complex molecular structures.
Construction of Molecular Structure Feature Sequences: This model iteratively extracts features for each atom within the molecule over k iterations, obtaining the features within each molecule’s k-neighborhood. These final molecular structure feature sequences are used as inputs to the Transformer, providing a solid foundation for subsequent solubility predictions. This process effectively captures complex interactions and long-range dependencies within molecules, thereby enhancing the model’s predictive accuracy.
Sequence Dependency Analysis with LSTM: In the YZS-Model, LSTM is used to handle the sequential data within molecules, particularly capturing long-term dependencies in molecular sequences. With its unique memory mechanism, LSTM addresses the gradient vanishing problem in traditional RNNs when processing long sequences. By integrating various aspects of molecular relationships, LSTM can further uncover temporal dependencies within molecular structure sequences, enhancing the model’s predictive capability. This allows the YZS-Model to excel in handling complex molecular sequence data, enabling more accurate predictions of drug solubility.
@@ 2.2 Data Preprocessing
In our study, we employed the RDKit library to convert SMILES (Simplified Molecular Input Line Entry System) strings into molecular objects. These molecular objects serve as the foundational data structure for further extraction and analysis of molecular features. RDKit is a widely used cheminformatics toolkit that allows for the processing and manipulation of chemical information, and it was essential for our feature extraction pipeline.
To accurately represent the intricate details of each molecule, we constructed a feature tensor that stores comprehensive information about the core attributes of each molecule. These attributes include:
Type of Atoms: The elements present in the molecule, such as carbon (C), nitrogen (N), oxygen (O), etc. Each element is encoded as a binary feature where 0 indicates the absence and 1 indicates the presence of that particular element. For a more detailed representation, each atom type is represented in a 66-bit one-hot encoded vector, as outlined in Table 1.
Number of Bonds: The number of bonds between atoms, indicating the molecule’s connectivity. This feature captures the total number of bonds each atom forms, represented in an 8-bit binary encoding.
Formal Charges: The formal charge on each atom, represented as a single integer value indicating the charge.
Free Radical Electrons: The number of unpaired electrons on each atom, represented as a single integer.
Hybridization Types: The hybridization state of each atom, such as sp, sp2, sp3, etc. This is encoded using a 7-bit one-hot encoding.
Aromaticity: Whether the atom is part of an aromatic system, encoded as a binary feature (true/false).
Number of Directly Bonded Hydrogen Atoms: This feature indicates the number of hydrogen atoms directly bonded to each atom, encoded using a 5-bit one-hot encoding.
Chirality and Stereochemistry: Information about the chirality (whether the atom is chiral) and the type of chirality (R or S configuration), with chirality being a binary feature and chirality type being a 2-bit one-hot encoding.
The complete set of atomic features utilized in our molecular object analysis is detailed in Table 1.
@@
For capturing the internal bond information within the molecules, we used two key data structures: ‘edge_attr’ and ‘edge_index’.
‘edge_attr’: This stores attributes related to the bonds, such as the type of bond (single, double, triple, aromatic), whether the bond is conjugated, its aromaticity, and whether the bond is part of a ring. Each bond type is encoded in a 4-bit binary feature as shown in Table 2.
‘edge_index’: This records the connectivity between atoms, including atom indices and types of bonds, providing a detailed structural framework of the molecule. The bond features are summarized in Table 2.
@@
These detailed features are subsequently integrated into a ‘torch_geometric’ Data object, which includes the node features, edge connectivity, and their attributes, forming a complete graph data structure for analysis and model training.
To visualize the distribution of these molecular features, we will present the figure 2 that depicts the frequency and diversity of the key attributes within our dataset. This figure will help illustrate the richness of the data and the broad range of molecular characteristics captured by our preprocessing pipeline.
Figure 2: Distribution of Molecular Features in the Training Dataset.
This visualization provides insight into the variability and complexity of the molecular structures analyzed in our study, highlighting the diversity of atomic and bond features present in the dataset.
@@
In our research, we adopted the version of the Cui2020 dataset segmented by Ahmad et al. for the training and testing phases of our machine learning model[21, 13]. The team led by Ahmad et al. randomly split the original Cui2020 dataset into ten equal and independent subsets, each termed a “fold”. To evaluate our model’s efficacy, we implemented a 10-fold cross-validation technique: during each cycle, the model trained on nine folds and was tested on the remaining one. This procedure was replicated ten times, ensuring that each fold was used as a test set once.
@@ 2.9 YZS-Model
The initial step of the model involves utilizing a Graph Convolutional Network (GCN) to process the graphical representation of drug molecules, effectively revealing the spatial structural features of the molecules. By analyzing the interactions between the atoms and their chemical bonds, the GCN layer can abstract complex feature vectors that describe the molecular topology, laying a solid foundation for in-depth structural analysis. After converting SMILES to a graph representation, where the nodes represent the feature vectors of each atom and the edges represent the bonds, the model uses a GCN to process this graphical representation of drug molecules. This effectively reveals the spatial structural features of the molecules. By analyzing the interactions between the atoms and their chemical bonds, the GCN layer abstracts complex feature vectors that describe the molecular topology, providing a robust basis for detailed structural analysis.
The model iteratively extracts features for each atom within the molecule over k iterations, obtaining the features within each molecule’s k-neighborhood. These final molecular structure feature sequences are used as inputs to the Transformer, providing a solid foundation for subsequent solubility predictions. This process effectively captures complex interactions and long-range dependencies within molecules, thereby enhancing the model’s predictive accuracy.
Subsequently, the model employs a Transformer architecture to further refine the feature output from the GCN layer. Utilizing self-attention mechanisms, the Transformer can discern the broad dependencies among atoms within the molecule, even if they are spatially distant in the molecular structure. This stage of analysis not only focuses on the microstructural details of the molecule but also captures the comprehensive internal interactions, forming a sophisticated feature representation that encompasses extensive global information. Atoms within a molecule can interact even if they are far apart, which can further influence solubility. Therefore, the Transformer architecture is used to analyze molecular features from a global perspective. By leveraging self-attention mechanisms, the Transformer can identify broad dependencies among atoms within the molecule, even if they are spatially distant in the molecular structure. This stage of analysis not only focuses on the microstructural details but also captures comprehensive internal interactions, forming a sophisticated feature representation that encompasses extensive global information.
Furthermore, the LSTM layer is used to deepen the understanding of the features generated by the Transformer architecture. Given the potential temporal dependencies inherent in molecular structure sequences, the LSTM effectively captures the temporal information in the sequence data through its unique gating mechanisms. This processing stage is crucial for revealing the dynamic nature of molecular features across sequences, providing deep temporal insights for solubility prediction.
Finally, through a fully connected layer, the model transforms the output from the LSTM layer into the final solubility prediction. These layers not only synthesize high-dimensional features from the GCN, Transformer, and LSTM but also link these features to the solubility of drug molecules through nonlinear transformations. The entire model is designed to effectively map the spatial and temporal characteristics of molecules to their physicochemical property of solubility, ensuring high accuracy and reliability in the solubility predictions for organic drugs.